OpenStudy (anonymous):

A right triangle has one leg that is 13 cm long. The hypotenuse is 27 cm long l. Find the area of the triangle to the nearest tenth .

OpenStudy (anonymous):

use a^2 + b^2 = c^2 do you see how 'a' and 'c' relate to your numbers on the triangles drawn? 

OpenStudy (anonymous):

Well yes . But how wpuld I plug the numbers in .? And what would I do next ?

OpenStudy (anonymous):

Thanks, and so after that i would square everythign and subtract from both sides right?

OpenStudy (anonymous):

so you should have: 27^2 - 13^2 = b^2, solve for b once you have b, you can solve for the area using: Area = (1/2)base*height

OpenStudy (anonymous):

i found what b= but i still am stuck on what to do next . :(

OpenStudy (anonymous):

ok, so now you have all the sides of the triangle, we next will find the area. from the numbers you have, how 'tall' is the triangle? and how 'wide' is the bottom side of the triangle?

OpenStudy (anonymous):

the triangle has one leg that is 13 cm long. the hypotenuse is 27 cm long. find the ara og the trianle to the nearest tenth . So far i had did : a^2=b^2=c^2 27^2-13^2=b^2 and i ended up with 560=b^2 so now i dont know whta to do next

OpenStudy (anonymous):

560=b^2 now take the square root of both sides so we get rid of the power of 2 on the b